If nC12 = nC8, find out what n is.
Found 2 solutions by richard1234, Edwin McCravy:
Answer by richard1234(7193)

n = 20

Note: nCk = nC(n-k)

Answer by Edwin McCravy(20055)

Use the fact that

 nCr = nC(n-r)

nC12 = nC8

Comparing those, you get the two equations:

r=12
n-r=8

Substitute 12 for r in the second one

n-12=8
   n=20

Checking with a TI-84 calculator:

nC12 = nC8

20C12 = 20C8

125970 = 125970

Edwin